Solve the system of equations by graphing.
x+y=8
x-y=0

Answer by mananth(16949)

x+y=8............1
x-y=0............2
equation1
when x= 0 y= 8 ( 0 , 8 )
when y = 0 x= 8 ( 8 , 0 ) 8
Equation 2
when x= 0 y= 0 ( 0 , 0 )
when y = 0 x= 0 ( 0 , 0 )

m.ananth@hotmail.ca